# MedTransformer: Accurate Alzheimer’s Disease Diagnosis for 3D MRI Images through 2D Vision Transformers

### 12 Jan 2024 v1

@@ 3.3 Fusion Attention Mechanism
To better combine the information among different slices and different views, we propose a simple but novel cross-attention mechanism. We call this mechanism fusion attention. The fusion attention mechanism tries to add the embeddings together directly. However, different from simply adding them together one by one, it adds the embeddings representing the patches but not the tokens. Please note that the $[class]$ token of each embedding has aggregated the information from one slice in previous encoders, so this operation aims to let the embeddings more focus on themselves when learning attention. At the same time, it can also extract the feature information from other slices or dimensions. Here also takes IntraCAE as an example:
@@
Let’s consider it using a more mathematical way with the example of two embeddings. The traditional attention mechanism is shown as Eq. 14. After fusing these two embeddings, the K matrix of the first embedding will consist of the K value corresponding to the $[class]$ token from the first embedding, and the K matrix corresponding to fusion embedding. The Q matrix is similar. After the matrix calculation, Eq. 16 greatly fuses the information from two embeddings while keeping some unique information from the special $[class]$ token.
@@
$QK^{T}=\begin{bmatrix}Q_{class_{1}}K_{class_{1}}&(Q_{1}+Q_{2})K_{class_{1}}\\ Q_{class_{1}}(K_{1}+K_{2})&(Q_{1}+Q_{2})(K_{1}+K_{2})\end{bmatrix}$ (16)
@@ 3.4 Morphology Augmentation
After visualizing the AD and normal images, we uncover a crucial insight: the atrophy of the brain mass size between these two kinds of images are significantly different. Based on this finding, we propose morphology augmentation, an augmentation method which can help to expand and reduce the size of the atrophy, causing the improvement of the model in distinguishing the two classes. This augmentation is based on scale erosion and dilation shown in Eq.  1718. f is the input image, $b_{N}$ is the erosion or dilation element, (x,y) and (s,t) are the coordinates in f and $b_{N}$ respectively.
@@
The visualization of morphology augmentation is shown in Fig. 3. We do erosion-based morphology augmentation to the Alzheimer’s Disease(AD) image and dilation-based morphology augmentation to the Normal Control(NC) image. The size of the atrophy is increased in AD but decreased in NC. With the morphology augmentation, we can not only augment the AD and NC images but also utilize the MCI data more efficiently. Because the MCI is the prodromal stage of AD, so we can do erosion-based morphology augmentation to the MCI data and classify them as AD, while doing dilation-based morphology augmentation to the MCI data and then classify them as normal data.
Figure 3: The Alzheimer’s Disease (AD) image, Normal Control (NC) image, and Mild Cognitive Impairment (MCI) image before (left) and after (right) morphology augmentation.
@@ 4 Experiments
Model name ADNI AIBL MIRIAD OASIS
@@
MedicalNet-18 0.827 0.793 0.856 0.847 0.793
MedicalNet-18 0.756 0.743 0.820 0.800 0.809
MedicalNet-34 0.571 0.669 0.847 0.782 0.585
MedicalNet-50 0.588 0.471 0.303 0.706 0.317
MedicalNet-101 0.558 0.425 0.320 0.656 0.291
MedicalNet-152 0.564 0.381 0.175 0.662 0.239
MedicalNet-200 0.556 0.361 0.125 0.665 0.220
COVID-ViT 0.506 0.573 0.644 0.338 0.715
Uni4Eye 0.622 0.646 0.697 0.647 0.688
MedTransformer 0.781 0.810 0.863 0.756 0.811
@@ 4.1 Experimental Dataset
To verify the efficiency of our MedTransformer, we use the dataset from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) for the training process. This dataset consists of MRI images of T1-weighted magnetic resonance imaging subjects. There are a total of 3,891 3D MR images in the dataset, including 1,216 normal cases (NC), 1,110 AD cases and 1,565 MCI cases. During the training, 878 normal images, 884 AD images and 1565 MCI images were split into the training set, with 72 normal images and 81 AD images as a validation set, together with 266 normal images and 145 AD images as a testing set.
At the same time, to evaluate the performance of our MedTransformer and other deep learning baseline models, we also consider other datasets as test sets. We mainly acquire them from three other institutions with the ADNI test dataset: Australian Imaging, Biomarker and Lifestyle Flagship Study of Ageing (AIBL), Minimal Interval Resonance Imaging in Alzheimer’s Disease (MIRIAD), and The Open Access Series of Imaging Studies (OASIS). The AIBL dataset contains a total of 481 cases which consist of 363 NC, 50 AD cases and 68 MCI cases. Note that when testing, we drop all the MCI cases. So there are total 413 images with 363 NC and 50 AD in AIBL dataset. The MIRIAD dataset contains a total of 523 cases which consist of 177 NC and 346 AD cases. The OASIS dataset contains a total of 2157 cases which consist of 1692 NC and 465 AD cases. Each of the images for any dataset is a 3D grayscale image, meaning that only has 1 channel as input.
@@ 4.2 Implementation Details
We implement consistent data pre-processing techniques to normalize and standardize MR images sourced from a multi-institutional database. We first do data augmentation in the following steps. we have followed closely the recommended protocol from the medical community [51] to process the data. Firstly, we do bias field correction with N4ITK method [44]. Next, we register each image to the MNI space [15, 14] with the ICBM 2009c nonlinear symmetric template by performing a affine registration using the SyN algorithm [2] from ANTs [1]. At the same time, the registered images were further cropped to remove the background to improve the computational efficiency. These operations result in images of size 169 × 208 × 179, with 1 mm isotropic voxels. Intensity rescaling, which was performed based on the minimum and maximum values, denoted as MinMax, was also set to be optional to study its influence on the classification results. Finally, the deep QC system [16] is performed to check the quality of the linearly registered data. The software outputs a probability indicating how accurate the registration is. We excluded the scans with a probability lower than 0.5. Overall, the registration process we perform on the data maps different sets of images into a single coordinate system to prepare the data for our later usage.
@@
We also use the Torchio library [36] to do data augmentation. We only apply this augmentation to the training dataset while keeping the validation and testing dataset unchanged. At the same time, we resize all the MR images with Scipy library [46] into 224$\times$224$\times$224 to better fit the input of our MedTransformer. Finally, we employed the zero-mean unit-variance method to normalize the intensity of all voxels within the images.
@@
For the training dataset, we leverage MCI data more smartly. We do morphology augmentation to the same MCI data, classify the MCI into NC after doing dilation-based augmentation, and classify it into AD after doing erosion-based augmentation. In this way, each MCI is used twice, significantly enlarging the dataset. At the same time, we also do morphology augmentation to AD and NC images randomly, with a probability of 0.5.
@@
After preprocessing the 3D MR images, we cut them into 2D slices along sagittal, coronal and axial views. Then we choose a number of slices in each view and concatenate them into a sequence. We choose 20 equidistant slices on each view and embed them into patch embedding similar to ViT. Here we choose the embed layer from  [43]. Then we use a total of 12 standard transformer attention layers, and 3 layers in each encoder, with 4 heads. At last, because we have three $[class]$ tokens, each representing a special view dimension, we use a classification MLP head, with input feature number 3$\times$256 and output feature number 2, aiming to figure out whether the image is from a disease or not.
@@
We implemented MedTransformer using a Pytorch library [35]. MedTransformer was trained using an AdamW optimizer with learning rate 0.00005. All other parameters are default. At the same time, we also took the advantage of cosine learning rate from [30]. Because this is a binary classification task, we used cross-entropy loss [56]. The training process used five 24G NVIDIA A5000 GPUs. Due to the memory capacity, we use 2 batches on each GPU, meaning a total batch size of 10.
@@ 4.3 Evaluation Between Baselines
Our MedTransformer was compared with various baseline models, including 3D CNN-based models: MedicalNet [8] and 3D transformer-based models: COVID-VIT [18], Uni4Eye [4]. There are various versions of MedicalNet, each of which is based on a basic Resnet [24] model, such that MedicalNet-10 is based on Resnet-10 respectively.
In this experiment, we chose a total of 90 slices as input, meaning 30 equidistant slices on each view. Because we found that the central part of the 3D images would be more important and consist of more useful information, we applied the important sampling method in our slice-picking stage. To be more specific, for a 224$\times$224×2absent2$\times$ 2× 224 image, we pick equidistant slices from $\rm 52^{nd}$ to $\rm 172^{nd}$ on each view.
The quantitative performance is presented in Table 1. We chose the model with the best validation accuracy on ADNI and then tested it on various Alzheimer’s disease datasets. Overall, MedTransformer achieves the best performance on i.i.d testing scenario (ADNI) as well as two of the three out-of-domain testing scenario (AIBL and OASIS). We believe these results show that MedTransformer is not only superior in Alzheimer’s diagnosis in i.i.d setting, but also fairly robust when the testing data is collected from different facilities. For MIRIAD dataset, MedTransformer also achieves a considerable result although not the best-performing one. We conjecture this is due to the disparity of data characteristics between ADNI and MIRIAD. In particular, we notice that MIRIAD has a distinct ratio of AD vs. normal control in the dataset in comparison to others.
@@ 4.4 Ablation Study
Slice Number ADNI AIBL MIRIAD OASIS
Val acc. Test acc. Test acc. Test acc. Test acc.
15 0.679 0.690 0.778 0.536 0.781
20 0.638 0.738 0.846 0.402 0.777
45 0.669 0.721 0.797 0.561 0.771
Ours (30) 0.781 0.810 0.863 0.756 0.811
Table 2: Comparison of accuracy between MedTransformer and three variants ablating with different numbers of 2D slices on each axial, coronal, and sagittal dimension in the four datasets.
@@
1+1+2+2 0.713 0.776 0.800 0.685 0.793
2+2+1+1 0.708 0.719 0.769 0.674 0.744
4+4+4+4 0.523 0.670 0.814 0.355 0.789
Ours (3+3+3+3) 0.781 0.810 0.863 0.756 0.811
Table 3: Comparison of accuracy between MedTransformer and three variants ablating with different numbers of transformer layers in each encoder in the four datasets.
@@
No Cross-Attension 0.571 0.683 0.838 0.485 0.785
Class Token Cross-Attention 0.656 0.637 0.654 0.710 0.606
Easy Concat Cross-Attention 0.622 0.655 0.719 0.653 0.681
Ours (Fusion Attention) 0.781 0.810 0.863 0.756 0.811
Table 4: Comparison of accuracy between MedTransformer and three variants ablating different cross attention mechanisms in the four datasets.
@@
w/o Morphology Augmentation 0.727 0.714 0.731 0.410 0.698
w/o Guide Eembedding 0.75 0.717 0.814 0.602 0.777
w/o Torchio 0.682 0.710 0.796 0.521 0.781
w/o Pretrained Weights 0.644 0.644 0.702 0.714 0.650
w/o Important Sampling 0.544 0.603 0.508 0.531 0.719
MedTransformer 0.781 0.810 0.863 0.756 0.811
Table 5: Comparison of accuracy between MedTransformer and five variants ablating different training augmentation settings in the four datasets.
To evaluate how effective each block is, we compared our MedTransformer with other variants, changing one setting each time. We first compare how different numbers of slices on each dimension will affect the final result. We chose numbers 15, 20, and 45 as comparisons, and our MedTransformer chooses 30 slices on each view dimension. Table 2 shows that fewer slices may not be enough for transformer encoders to extract all the useful information. However, more slices may lead to overfitting problems due to noisy information extracted.
Then we also changed the transformer attention layers of each encoder. ViT uses 12 layers to embed the sequence, so we also tried to cut 12 layers into 4 encoder blocks equally, meaning 3 layers for each encoder block. At the same time, we also investigate into how the number of layers will affect our MedTransformer performance. The results are shown in Table 3, there are four numbers in each variant, each one corresponding to an encoder block. Such as 1+1+2+2 meaning that the shared self-attention encoders, dimension-specific self-attention encoders, intra-dimension cross-attention encoders and inter-dimension cross-attention encoders have 1, 1, 2, 2 transformer attention layer respectively. The result shows that MedTransformer outperforms all the variants on test accuracy in four different datasets, meaning that 3 layers for each block has its superiority.
Table 4 shows how different cross-attention mechanisms will affect the final result. The first variant: No Cross-Attention, meaning that we didn’t apply any cross-attention mechanism in the last two encoder blocks. Class Token Cross-Attention is a variant of Eq. 15. It adds the $[class]$ token embedding up but not the embedding behind the $[class]$ token. For the easy concat cross-attention mechanism, it simply concatenates the embeddings from different slices and view dimensions into a whole large embedding. Our proposed Fusion Attention achieves more than 12% improvements to the ADNI test result while demonstrating superiority on other testing datasets, verifying that fusion attention cannot only fuse the information while keeping the unique information in each embedding.
Table 5 shows other variables in our settings. We delete one important setting in each variant to see the results. MedTransformer outperforms all variant models in all four datasets by 9.03%, 4.3%, 4.2% and 2.8%, respectively. The results show the great capability of different settings in augmenting the model learning ability to classify 3D MRI.

### 29 Feb 2024 v2

@@ 3.2 Fusion Attention Mechanism
The above architecture will allow us to learn the intricies of AD pathologies along three different dimensions. However, the complicatedness of AD will require the model to thoroughly integrate the information from these three dimensions. Thus, we propose a cross-attention mechanism, namely funsion attention. The fusion attention adds the embeddings together directly. However, different from simply adding them together one by one, it adds the embeddings representing the patches but not the tokens. Note that the $[class]$ token of each embedding has aggregated the information from one slice in previous encoders, so this operation will let the embeddings more focus on themselves when learning attention. At the same time, it can also extract the feature information from other slices or dimensions. The fusion attention applied to both IntraCAE and InterCAE, but here we use IntraCAE as an example:
@@
In a more formal way, the traditional attention mechanism is shown as Eq. 9. After fusing these two embeddings, the $K$ matrix of the first embedding will consist of the $K$ value corresponding to the $[class]$ token from the first embedding, and the $K$ matrix corresponding to fusion embedding, similarly for $Q$ matrix. After the matrix calculation, Eq. 11 fuses the information from two embeddings while keeping some unique information from the special $[class]$ token.
@@
$Q_{1}K_{1}^{T}=\begin{bmatrix}Q_{class_{1}}K_{class_{1}}&(Q_{1}+Q_{2})K_{class_{1}}\\ Q_{class_{1}}(K_{1}+K_{2})&(Q_{1}+Q_{2})(K_{1}+K_{2})\end{bmatrix}$ (11)
@@ 3.3 Morphology Augmentation
A key characteristic of the AD-plagued brain is that, as the disease progresses, an increasing amount of brain mass will suffer from atropy. When this process is reflected in brain imaging, the there will be empty “holes” of the brain if one has AD. Based on this, we propose a morphology augmentation, an augmentation method which help to expand and reduce the size of the atrophy, causing the improvement of the model. This augmentation is based on atrophy expansion and atrophy reduction shown in Eq.  1213. $f$ is the input image, $b_{N}$ is the atrophy expansion or atrophy reduction element, $(x,y)$ and $(s,t)$ are the coordinates in $f$ and $b_{N}$ respectively.
@@
We apply atropy expansion augmentation to AD images and MCI images and label the resultant images as AD; on the other hand, we apply atropy reduction augmentation to Normal Control(NC) images and MCI images and label the resultant images as NC, where MCI is the prodromal stage of AD. The visualization of morphology augmentation is shown in Fig. 3.
Figure 3: The visualization of Alzheimer’s Disease (AD) image, Normal Control (NC) image and Mild Cognitive Impairment (MCI) image. The left is the raw image and the right is the augmented image. The cerebral ventricle (red circle) has a significant difference in size for AD and NC.
3.4 Adaptive Training Strategy
To further investigate the potential of ADAPT, we propose an attention score based training strategy in order to allow our model to extract more features from the more important dimension with limited size of inputs. We calculate the attention score of each dimension after the final inter-dimension cross attention encoder layer according to Eq. 14. Because our $[class]$ token is dimension specific, so we just calculate the attention score of the $[class]$ token as the representation of the special dimension. This strategy allows our network to adaptively choose the slice number of each dimension while updating itself.
$H_{dim}=softmax(\frac{Q_{class_{dim}}K^{T}}{\sqrt{d_{k}}})V$ (14)
0:  3D MRI Training set $T$, initial slice number list $\psi$, model ADAPT $\Theta$, total slice number $n_{total}$
0:  Updated model $\Theta$, final list $\psi$
1:  while $Training$ do
2:     Sample 2D data $\delta$ with $T$ and $\psi$
3:     Train($\Theta$,$\delta$)
4:     Calculate score using Eq. 14 for three dimensions
5:     Calculate cross-entropy loss and update $\Theta$
6:     if p then
7:        Update $\psi$ according to Eq 1615.
8:     else
9:        Initialize($\psi$)
10:     end if
11:  end while
Algorithm 1 ADAPT Training Strategy
We then adaptively update the slice number of each dimension based on normalized attention scores using Eq. 1516, where $n$ is the total slice number and $\psi$ is the slice number list. Here we also constrain the selection pool to make sure the model will attend across multiple attentions. The full training strategy is shown in algorithm 1. To avoid the model will stick with certain view dimension after the first choice, we also allow the model to change the attention with certain probabilities p.
$n_{dim}=round(\frac{\hat{H_{dim}}}{\sum_{r\in\textsc{$\psi$}}r}*n_{total})$ (15)
$n_{dim}=\begin{cases}n_{min},n_{dim}\geq n_{min}\\ n_{max},n_{dim}\leq n_{max}\end{cases}$ (16)
@@ 4 Experiments
Model name Model size (#params) GFLOPs ADNI AIBL MIRIAD OASIS
@@
MedicalNet-10 (Chen et al., 2019) 17,723,458 225.7 0.827 0.843 0.856 0.847 0.793
MedicalNet-18 (Chen et al., 2019) 36,527,938 492.6 0.756 0.793 0.820 0.800 0.809
MedicalNet-34 (Chen et al., 2019) 66,837,570 910.8 0.571 0.669 0.847 0.782 0.585
MedicalNet-50 (Chen et al., 2019) 59,626,818 666.8 0.588 0.471 0.303 0.706 0.317
MedicalNet-101 (Chen et al., 2019) 98,672,962 1181.1 0.558 0.425 0.320 0.656 0.291
MedicalNet-152 (Chen et al., 2019) 130,831,682 1604.8 0.564 0.381 0.175 0.662 0.239
3D ResNet-34 (He et al., 2016) 63,470,658 341.1 0.506 0.618 0.779 0.453 0.745
3D ResNet-50 (He et al., 2016) 46,159,170 256.9 0.519 0.618 0.776 0.657 0.720
3D ResNet-101 (He et al., 2016) 85,205,314 391.1 0.569 0.526 0.601 0.426 0.433
3D DenseNet-121 (Huang et al., 2017) 11,244,674 260.5 0.609 0.675 0.816 0.408 0.772
3D DenseNet-201 (Huang et al., 2017) 25,334,658 286.8 0.564 0.612 0.672 0.553 0.671
Knowledge4D (Zhou et al., 2023b) 33,162,880 633.9 0.603 0.728 0.845 0.658 0.784
I3d (Carreira & Zisserman, 2017) 12,247,332 191.0 0.513 0.629 0.674 0.448 0.607
FCNlinksCNN (Qiu et al., 2020) 310,488,372 375.6 0.423 0.614 0.708 0.370 0.691
COVID-ViT (Gao et al., 2021) 78,177,282 448.6 0.506 0.573 0.644 0.662 0.715
Uni4Eye (Cai et al., 2022) 340,324,866 78.4 0.622 0.646 0.697 0.647 0.688
ADAPT 9,695,490 46.3 0.904 0.924 0.910 0.903 0.817
@@ 4.1 Experimental Dataset
To verify the effectiveness of our ADAPT, we use the dataset from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) for the training process. This dataset consists of MRI images of T1-weighted magnetic resonance imaging subjects. There are a total of 3,891 3D MRI images in the dataset, including 1,216 normal cases (NC), 1,110 AD cases and 1,565 MCI cases. During the training, 878 normal images, 884 AD images and 1565 MCI images were split into the training set, with 72 normal images and 81 AD images as a validation set, together with 266 normal images and 145 AD images as a testing set.
At the same time, to evaluate the performance of our ADAPT and other deep learning baseline models, we also consider other datasets as test sets. We mainly acquire them from three other institutions with the ADNI test dataset: Australian Imaging, Biomarker and Lifestyle Flagship Study of Ageing (AIBL), Minimal Interval Resonance Imaging in Alzheimer’s Disease (MIRIAD), and The Open Access Series of Imaging Studies (OASIS). The AIBL dataset contains a total of 413 images with 363 NC and 50 AD after dropping all MCI cases. The MIRIAD dataset contains a total of 523 cases which consist of 177 NC and 346 AD cases. The OASIS dataset contains a total of 2157 cases which consist of 1692 NC and 465 AD cases. Each of the images for any dataset is a 3D grayscale image.
@@ 4.2 Implementation Details
We implement consistent data pre-processing techniques to normalize and standardize MRI images sourced from a multi-institutional database. We first do data augmentation in the following steps. we have followed closely the recommended protocol from the medical community (Wen et al., 2020) to process the data. Firstly, we do bias field correction with N4ITK method (Tustison et al., 2010). Next, we register each image to the MNI space (Fonov et al., 2009, 2011) with the ICBM 2009c nonlinear symmetric template by performing a affine registration using the SyN algorithm (Avants et al., 2014) from ANTs (Avants et al., 2008). At the same time, the registered images were further cropped to remove the background to improve the computational efficiency. These operations result in 1 mm isotropic voxels for each image. Intensity rescaling, which was performed based on the minimum and maximum values, denoted as MinMax, was also set to be optional to study its influence on the classification results. Finally, the deep QC system (Fonov et al., 2018) is performed to check the quality of the linearly registered data. The software outputs a probability indicating how accurate the registration is. We excluded the scans with a probability lower than 0.5. Overall, the registration process we perform on the data maps different sets of images into a single coordinate system to prepare the data for our later usage.
@@
We also use the Torchio library (Pérez-García et al., 2021) in the training set. At the same time, we resize all the MRI images with Scipy library (Virtanen et al., 2020) into 224$\times$224$\times$224 to better fit the input of our ADAPT. Finally, we employed the zero-mean unit-variance method to normalize the intensity of all voxels within the images.
@@
For the training dataset, we apply morphology augmentation to the same MCI data, classify the MCI into NC after doing atrophy reduction augmentation, and classify it into AD after doing atrophy expansion augmentation. In this way, each MCI is used twice, significantly enlarging the dataset. At the same time, we also do morphology augmentation to AD and NC images randomly, with a probability of 0.5.
@@
After preprocessing the 3D MRI images, we cut them into 2D slices along sagittal, coronal and axial views. Then we choose 16 slices in each view as the initial data and concatenate them into a sequence. We choose equidistant slices on each view and embed them into patch embedding similar to ViT. Here we choose the embed layer from  (Touvron et al., 2022). Then we use a total of 6 standard transformer attention layers, and 1 layer for each of the first two encoders, 2 layers for each of the last two encoders, with 4 heads. For the adaptive training strategy, we set the probability p as 0.8. At last, because we have three $[class]$ tokens, each representing a special view dimension, we use a classification MLP head, with input feature number 3$\times$256 and output feature number 2, aiming to figure out whether the image is from a disease or not.
@@
We implemented ADAPT using a Pytorch library (Paszke et al., 2019). ADAPT was trained using an AdamW optimizer with a learning rate of 0.00005. All other parameters are default. At the same time, we also took the advantage of cosine learning rate from (Loshchilov & Hutter, 2016). We treat this as a binary classification task, so we use cross-entropy loss (Zhang & Sabuncu, 2018). The training process used 2 80G NVIDIA A800 GPUs. Due to the memory capacity, we use 6 batches on each GPU, meaning a total batch size of 12.
@@ 4.3 Evaluation Between Baselines
Our ADAPT was compared with various baseline models, including 3D CNN-based models: 3D DenseNet (121, 201) (Huang et al., 2017), 3D ResNet (34, 50, 101) (He et al., 2016) because they have been widely used for AD classification (Korolev et al., 2017; Ruiz et al., 2020; Yang et al., 2018; Zhang et al., 2021). We also add other baselines to show the capability of our ADAPT, including: MedicalNet (Chen et al., 2019), I3D (Carreira & Zisserman, 2017), FCNlinksCNN (Qiu et al., 2020) and Knowledge4D (Zhou et al., 2023b). There are various versions of MedicalNet, each of which is based on a basic Resnet (He et al., 2016) model, such that MedicalNet-10 is based on Resnet-10 respectively. We also compare our method with 3D transformer-based models: COVID-VIT (Gao et al., 2021), Uni4Eye (Cai et al., 2022).
In the experiment, we chose a total of 48 slices as input, meaning 16 equidistant slices on each view as initial. Because we found that the central part of the 3D images would be more important and consist of more useful information, we applied the important sampling method in our slice-picking stage. To be more specific, for a 224$\times$224×2absent2$\times$ 2× 224 image, we pick equidistant slices from $\rm 52^{nd}$ to $\rm 172^{nd}$ on each view.
The quantitative performance is presented in Table 1. We chose the model with the best validation accuracy on ADNI and then tested it on various Alzheimer’s disease datasets. We also recorded the total parameters and GFLOPs of each model. Overall, ADAPT achieves the best performance on i.i.d testing scenario (ADNI) as well as all out-of-domain testing scenarios (AIBL, MIRIAD and OASIS). We believe these results show that ADAPT is not only superior in Alzheimer’s diagnosis in i.i.d setting, but also fairly robust when the testing data is collected from different facilities. At the same time, our model has the least parameters and GLOPs, demonstrating the success of our novel method in attacking the AD diagnosing task using 2D based model.
The best performance is achieved when ADAPT chooses 10 slices from saggital view, and 19 slices from the coronal and axial view respectively. As compared with table 4, we found the interesting facts that coronal and axial view may contain more differential relationships about cortex and ventricle of AD and NC, which can help the model learn the special attention features accurately.
@@ 4.4 Ablation Study
1+1+1+1 0.713 0.776 0.800 0.685 0.793
2+2+1+1 0.770 0.811 0.863 0.903 0.716
2+2+2+2 0.881 0.911 0.897 0.669 0.800
3+3+3+3 0.917 0.895 0.907 0.723 0.806
Ours (1+1+2+2) 0.904 0.924 0.910 0.903 0.817
Table 2: Comparison of accuracy between ADAPT and three variants ablating with different numbers of transformer layers in each encoder in the four datasets.
@@
No Cross-Attention 0.719 0.627 0.810 0.710 0.675
Class Token Cross-Attention 0.878 0.848 0.864 0.709 0.606
Easy Concat Cross-Attention 0.917 0.783 0.723 0.806 0.681
Ours (Fusion Attention) 0.904 0.924 0.910 0.903 0.817
Table 3: Comparison of accuracy between ADAPT and three variants ablating different cross attention mechanisms in the four datasets.
@@
w/o Morphology Augmentation 0.604 0.738 0.889 0.828 0.792
w/o Adaptive Training 0.788 0.855 0.883 0.902 0.807
w/o Guide Eembedding 0.750 0.800 0.863 0.869 0.813
w/o Torchio 0.678 0.755 0.825 0.664 0.783
w/o Pretrained Weights 0.731 0.738 0.877 0.746 0.800
w/o Important Sampling 0.713 0.712 0.757 0.767 0.787
ADAPT 0.904 0.924 0.910 0.903 0.817
Table 4: Comparison of accuracy between ADAPT and five variants ablating different training augmentation settings in the four datasets.
To evaluate how effective each block is, we compared our ADAPT with other variants, changing one setting each time. We first changed the transformer attention layers of each encoder. We investigate how the number of layers will affect our ADAPT performance. The results are shown in Table 2, there are four numbers in each variant, each one corresponding to an encoder block. Such as 1+1+2+2 meaning that the shared self-attention encoders, dimension-specific self-attention encoders, intra-dimension cross-attention encoders and inter-dimension cross-attention encoders have 1, 1, 2, 2 transformer attention layer respectively. The result shows that ADAPT outperforms all the variants on test accuracy in three out of four datasets (ADNI, AIBL, MIRIAD) while using small enough memory.
Table 3 shows how different cross-attention mechanisms will affect the final result. The first variant: No Cross-Attention, meaning that we didn’t apply any cross-attention mechanism in the last two encoder blocks. Class Token Cross-Attention is a variant of Eq. 10. It adds the $[class]$ token embedding up but not the embedding behind the $[class]$ token. For the easy concat cross-attention mechanism, it simply concatenates the embeddings from different slices and view dimensions into a whole large embedding. Our proposed Fusion Attention achieves more than 7% improvements to the ADNI test result while demonstrating superiority on other testing datasets, verifying that fusion attention cannot only fuse the information while keeping the unique information in each embedding.
Table 4 shows other variables in our settings. We delete one important setting in each variant to see the results. ADAPT outperforms all variant models in all four datasets by 4.9%, 2.1%, 0.1% and 1.0%, respectively. The results show the great capability of different settings in augmenting the model learning ability to classify 3D MRI.